OpenStudy (anonymous):

how do I subtract 1 from 5/2?

OpenStudy (aravindg):

first make denominators same 1 is same as 2/2

OpenStudy (aravindg):

5/2-1=5/2-2/2=3/2

OpenStudy (aravindg):

Done :)

OpenStudy (anonymous):

why is 1 2/2?

OpenStudy (aravindg):

1 is same as 1/1 multiply numerator and denominator with 2 2/2

OpenStudy (aravindg):

got it?

OpenStudy (anonymous):

so since 2 and

OpenStudy (anonymous):

1 both have the denominator of 2 I multiple 2 to the top and bottom of 1 giving me 2/2?

OpenStudy (aravindg):

yes

OpenStudy (anonymous):

if instead of 1 it was a 5 then I would see what is the denominator in 2 and 5 which is 10 so my answer would be 50/10 = 5/1?

OpenStudy (anonymous):

5/2 - 1 = 3/2 and the mixed fraction would be 1 and 1/2

OpenStudy (aravindg):

if it wa 5 instead of 1 u multiply with 2 10/2-5/2=5/2

OpenStudy (anonymous):

why 2?

OpenStudy (aravindg):

because we need to make denominators same

OpenStudy (aravindg):

the denominator of second term is 2 .So we need to make denominator of first term also 2

OpenStudy (anonymous):

ok that makes a little bit more sense

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

5/2 - 1 = 3/2 (3/2)

OpenStudy (kenljw):

In General 1/a +-1/b =(1/a)(b/b)+-(1/b)(a/a) = (b+-a)/ab Division is the inverse of Multiplication a/a = 1 any number divided by itself is 1 Multiplicative Identity 1 time a = a any number times one is itself